OpenStudy (anonymous):

I'm bad at fractions so I don't understand how to solve for X in this question. X- 3/4=6 2/3

OpenStudy (anonymous):

The first thing I would do is change 6 2/3 into a fraction. To do this, you multiply 6 by the denominator (3) and add 2: 6*3+2= 18+2 = 20 So you have x-3/4 = 20/3 With me so far?

OpenStudy (anonymous):

Yes, So I think after that you take 3/4 and add it to each side don't you?

OpenStudy (anonymous):

Yes, but you need to modify 3/4 before you can add it. To do that, we need a common denominator for the two fractions. The easy way to get that in this equation is to multiply them together -> 3*4=12 So, for 3/4, you multiply the top by 3 and get 9/12. For 20/3 you multiply the top by 4 and get 80/12. Make sense? Then you have: x-9/12=80/12

OpenStudy (anonymous):

I remember getting that far but not sure if I was doing it right, that makes sense so far.

OpenStudy (anonymous):

Okay. So now you can add 9/12 to both sides: x = 89/12 When you multiply the fractions to get a common denominator, you have to multiply by one. In this problem, we multiplied 3/4 * 3/3 to get 3*3/4*3. The other side multiplied 20/3 * 4/4 to get 20*4/3*4. This gives you the common denominator of twelve so that you can just add the top. Good luck!

OpenStudy (anonymous):

Thank you